A stone projected at an angle q with horizontal from the roof of a tall building falls on the ground after three seconds. Two seconds after the projected it was again at the level of projection. Then the height of the building is


        (a)    15 m


        (b)    5 m


        (c)    25 m


        (d)    20 m


Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

(a)

    TAC = 2 sec.


    So, (2u sinx)/g = 2 => u sinx = g = 10 m/s

 

    Now y = uyt + 1/2 ayt2 ==>– h = u sinx × 3 -1/2 g × 32

     –h = 10 × 3 -1/2 × 10 × 9 => – h = 30 – 45

     h = 45 – 30 = 15 m